# Beyond Accuracy: Evaluating the Reasoning Behavior of Large Language Models - A Survey

### v1

@@ 4.2 Rationale-Based Evaluation
In contrast to conclusion-based evaluation schemes, rationale-based evaluation methods focus on examining the reasoning trace generated by the model, typically assessing its logical validity and coherence. Given the variability in the model’s reasoning, a spectrum of assessment techniques exist. For rationales following a highly consistent format —either through structured contexts or methods that guide the model’s response style (Wan et al., 2024) — rationales can be parsed into more formalized representations such as first-order logic (Saparov & He, 2023), computation graphs (Dziri et al., 2023), or causal graphs (Willig et al., 2022), facilitating a more granular examination. Alternatively, interpretable quantitative metrics, such as ROSCOE (Golovneva et al., 2023) or RECEVAL (Prasad et al., 2023), may be utilized to evaluate the rationales’ semantic alignment with the reasoning task, their coherence, factual consistency, and logical validity. In instances where rationales are less structured, qualitative inspections are commonly employed, either through diagnostic agents (Huang et al., 2024), or human judgements (Mondorf & Plank, 2024).
@@ 4.3 Interactive Evaluation
Similar to the principle of dynamic assessment within psychology (Haywood & Tzuriel, 2002), interactive evaluation offers a framework to engage with LLMs during the evaluation. This allows for more flexible assessments, tailored to the model’s specific reasoning behavior. For instance, adaptive evaluations dynamically select reasoning tasks based on the model’s responses, thus providing deeper insights into its capabilities and limitations beyond what static questionnaires can offer (Zhuang et al., 2023). Dialectic evaluation methods assess the model’s reasoning in dialogue form, for example by challenging the model’s conclusions (Wang et al., 2023), or engaging it in game-theoretical scenarios (Bertolazzi et al., 2023). While interactive evaluations offer nuanced insights, they lack the structure of traditional methods, posing challenges in terms of standardization and reproducibility.
@@ 4.4 Mechanistic Evaluation
Mechanistic evaluations of LLMs delve into the underlying processes that drive the model’s responses, aiming to uncover the “how” and “why” within their reasoning process. By analyzing the internal mechanisms, such as attention patterns (Hou et al., 2023), activation flows (Dutta et al., 2024), and the functional attributes of individual layers (Pirozelli et al., 2023), deeper insights into the model’s operational logic can be gained. Focusing on the model’s intrinsic processes, this framework contrasts with previous approaches, drawing parallels to the study of human reasoning from a neuroscientific perspective (Papo, 2015).
@@ 5 Discussion
Despite the remarkable performance of LLMs in reasoning tasks (Bubeck et al., 2023; Fu et al., 2023), our review suggests that current models more closely resemble stochastic parrots (Bender et al., 2021) than systematic reasoners. As reviewed in Section 3, the models’ reasoning behavior reveals significant conceptual errors and limitations, particularly in scenarios not covered by their training data. This underscores a limited functional linguistic competence of LLMs, as highlighted by Mahowald et al. (2024). It is likely that the success of current LLMs in reasoning tasks may predominantly reflect their extensive training data, rather than genuine reasoning capabilities. This highlights the necessity for more nuanced analyses of the models’ reasoning behavior, especially in novel scenarios. While human reasoning is not infallible (Johnson-Laird, 2010), the human capacity for robust reasoning and generalization from limited data is unmatched by current LLMs. Various studies point to the fundamental differences between human reasoning and that of LLMs, especially the models’ restrictive autoregressive pre-training objective (McCoy et al., 2023; Shanahan, 2024; Lenci, 2023). We advocate for more research—especially on reasoning behavior—within both humans and LLMs, to better discern and comprehend the essential components missing in LLMs, which are crucial for robust and systematic reasoning.
@@ 6 Conclusion
This survey provides a comprehensive review of literature that evaluates LLM-based reasoning beyond mere task accuracy, offering deeper insights into the reasoning behavior of large language models. We discuss the behavior of LLMs across three core reasoning tasks (RQ1), assessing logical, mathematical and causal reasoning. Furthermore, we outline and categorize predominant evaluation frameworks (RQ2). Our findings indicate that although LLMs demonstrate proficiency in reasoning problems that align with their training data, they often encounter significant conceptual difficulties in out-of-distribution scenarios. Based on these insights, we caution against shallow accuracy metrics on static reasoning tasks and advocate for more sophisticated reasoning assessments as presented in Section 4 in the future.
@@ Appendix A Human-Like Reasoning Behavior in LLMs
Recent research has begun to explore the extent to which human reasoning behaviors are reflected in large language models, given their training on human-generated data. Various studies delve into the manifestation of human-like reasoning behaviors within LLMs in the context of deductive reasoning tasks. For instance, Eisape et al. (2023) compare the behaviors of humans and LLMs, particularly those from the PaLM 2 series, in syllogistic reasoning tasks. The findings reveal that, similar to humans, LLMs are prone to logical fallacies and cognitive biases such as ordering effects. This susceptibility persists across model sizes, though larger models tend to engage in more deliberate reasoning, showing a reduced sensitivity towards these errors. A notable finding indicates that LLMs, unlike humans, rarely produce the “nothing follows” response, even when it represents the accurate deduction. In a similar vein, Dasgupta et al. (2022) demonstrate that LLMs, akin to humans, exhibit content effects in predicate logic problems, i.e. their reasoning is influenced by the semantic content of the task. This effect is evident in various LLMs including ChatGPT, PaLM 2, and Chinchilla (Hoffmann et al., 2022). In particular, analyses indicate that the models’ reasoning in syllogisms is biased by the believability of the conclusion, a behavior known as belief bias (Klauer et al., 2000). Further research corroborates the influence of semantic content on the models’ reasoning. Ando et al. (2023) demonstrate a belief-bias in models like ChatGPT, RoBERTa and BART (Lewis et al., 2020) by comparing the models’ behavior on abstract, belief-consistent, and belief-inconsistent syllogisms. Their analyses also uncover a predisposition towards conversion errors and atmosphere effects, where logical quantifiers are either misinterpreted or lead to uninformed inferences, respectively (Tversky & Kahneman, 1974). Itzhak et al. (2023) further show that methods such as instruction tuning and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) may amplify cognitive biases within LLMs. Mondorf & Plank (2024) delve into the inferential strategies of open-access LLMs in problems of propositional logic. Comprehensive manual evaluations of the models’ rationales reveal that LLMs utilize inferential strategies similar to those employed by human reasoners. Their evaluations further underline difficulties of LLMs with logical negations and their vulnerability to logical fallacies commonly observed in human reasoning. Similarly, McKenzie et al. (2023) indicate that LLMs tend to replicate human-like logical errors when engaging with the logical principle of modus tollens, suggesting an imitation of flawed reasoning patterns from their training data. Notably, this trend becomes more pronounced with increasing model size, a phenomenon denoted as inverse scaling.
@@ Appendix B Further Details on Evaluation Methods
FAC2E Use capability-specific instructions to elicit intermediate structured reasoning steps (crystallized and fluid reasoning). Evaluate each step separately, using ground truth responses and the BARTScore-Recall metric. Wang et al. (2024)

### v2

@@ 4.2 Rationale-Based Evaluation
In contrast to conclusion-based evaluation schemes, rationale-based evaluation methods focus on examining the reasoning trace generated by the model, typically assessing its logical validity and coherence. While this approach allows for more nuanced insights into the model’s reasoning behavior, rationale-based evaluation schemes are often more challenging to automate and scale. Given the variability in the model’s reasoning, a spectrum of assessment techniques exist. For rationales following a highly consistent format—either through structured contexts or methods that guide the model’s response style (Wan et al., 2024)—rationales can be parsed into more formalized representations such as first-order logic (Saparov & He, 2023), computation graphs (Dziri et al., 2023), or causal graphs (Willig et al., 2022), facilitating a more granular examination. Alternatively, interpretable quantitative metrics, such as ROSCOE (Golovneva et al., 2023) or RECEVAL (Prasad et al., 2023), may be utilized to evaluate the rationales’ semantic alignment with the reasoning task, their coherence, factual consistency, and logical validity. In instances where rationales are less structured, qualitative inspections are commonly employed, either through diagnostic agents (Huang et al., 2024), or human judgments (Mondorf & Plank, 2024).
@@ 4.3 Interactive Evaluation
Similar to the principle of dynamic assessment within psychology (Haywood & Tzuriel, 2002), interactive evaluation offers a framework to engage with LLMs during the evaluation. This approach allows for more flexible assessments tailored to the model’s specific reasoning behavior. Although such evaluations are often costly and challenging to scale, several techniques have been developed to automate the process. For instance, adaptive evaluations dynamically select reasoning tasks based on the model’s responses, thus providing deeper insights into its capabilities and limitations beyond what static questionnaires can reveal (Zhuang et al., 2023). Dialectic evaluation methods assess the model’s reasoning in dialogue form, for example, by challenging the model’s conclusions (Wang et al., 2023), or engaging it in game-theoretical scenarios (Bertolazzi et al., 2023). While interactive evaluations yield nuanced insights, they lack the structure of traditional methods, posing challenges in terms of standardization and reproducibility.
@@ 4.4 Mechanistic Evaluation
Mechanistic evaluations of LLMs delve into the underlying processes that drive the model’s responses, aiming to uncover the “how” and “why” within their reasoning processes. By analyzing the internal mechanisms such as attention patterns (Hou et al., 2023), activation flows (Dutta et al., 2024), and the functional attributes of individual layers (Pirozelli et al., 2023), deeper insights into the model’s operational logic can be gained, as illustrated in Section 3.1.1. Focusing on the model’s intrinsic processes, this framework contrasts with previous approaches, drawing parallels to the study of human reasoning from a neuroscientific perspective (Papo, 2015). Nonetheless, current methods remain compute-intensive, and their findings may not always generalize across different models and tasks (Bereska & Gavves, 2024; Ferrando et al., 2024).
@@ 5 Discussion
Despite the notable performance of large language models in prominent reasoning tasks (Bubeck et al., 2023; Fu et al., 2023), our review suggests that current models more closely resemble stochastic parrots (Bender et al., 2021) than systematic reasoners. As discussed in Section 3, we find that although many LLMs demonstrate proficiency in reasoning problems that align with their training data, the models’ reasoning behavior reveals significant conceptual errors and limitations in out-of-distribution scenarios. As highlighted by Mahowald et al. (2024), this suggests a limited functional linguistic competence in LLMs. It is likely that the apparent success of LLMs in reasoning tasks predominantly reflects their ability to memorize the extensive data they have been trained on (Wu et al., 2024; Dziri et al., 2023). Recent studies indicate that a substantial amount of benchmark datasets has been leaked to current LLMs (Balloccu et al., 2024; Xu et al., 2024), raising concerns about the insights derived from their performance on such benchmarks. Therefore, we advocate for more nuanced analyses of the models’ reasoning behavior, particularly in novel scenarios that the models have not previously encountered.
While human reasoning is not infallible (Johnson-Laird, 2010), the human capacity for robust reasoning and generalization from limited data remains unmatched by current LLMs. Various studies point to the fundamental differences between human reasoning and that of LLMs, especially the models’ restrictive autoregressive pre-training objective (McCoy et al., 2023; Shanahan, 2024; Lenci, 2023). We call for further research—particularly on reasoning behavior—within both humans and LLMs to better discern and comprehend the essential components missing in LLMs, which are crucial for robust and systematic reasoning.
Evaluation Method Advantages Disadvantages
Conclusion-based evaluation Allows for controlled setups Limited insights
Provides metrics for comparison Less reliable
Easy to automate and scale
Easy to reproduce
Rationale-based evaluation Offers more nuanced insights Difficult to automate and scale
More robust in certain scenarios Might require expert interpretation
Interactive evaluation Highly flexible Expensive
Customizable to model behavior Difficult to automate and scale
Less standardized and reproducible
Mechanistic evaluation Identifies features or circuits responsible for specific behaviors Findings may not generalize across tasks or models
Supports direct interventions on model internals Results may be hard to interpret Compute-intensive
Table 1: Comparison of strengths and limitations of the different evaluation methods.
@@ 6 Conclusion
This survey provides a comprehensive review of literature that evaluates LLM-based reasoning beyond mere task accuracy, offering deeper insights into the reasoning behavior of large language models. We discuss the behavior of LLMs across three core reasoning tasks (RQ1), assessing logical, mathematical and causal reasoning. Furthermore, we outline predominant evaluation frameworks and compare their respective strengths and limitations (RQ2). Our findings indicate that although LLMs demonstrate proficiency in reasoning problems that align with their training data, they often encounter significant conceptual challenges in out-of-distribution scenarios. Considering these insights and the recent issue of benchmark dataset leakage (Balloccu et al., 2024; Xu et al., 2024), we caution against relying solely on shallow accuracy metrics for static reasoning tasks. Instead, we advocate for more sophisticated reasoning assessments, such as those discussed in Section 4.
@@ Appendix A Human-Like Reasoning Behavior in LLMs
Recent research has begun to explore the extent to which human reasoning behaviors are reflected in large language models, given their training on human-generated data. Various studies delve into the manifestation of human-like reasoning behaviors within LLMs in the context of deductive reasoning tasks. For instance, Eisape et al. (2024) compare the behaviors of humans and LLMs, particularly those from the PaLM 2 series, in syllogistic reasoning tasks. The findings reveal that, similar to humans, LLMs are prone to logical fallacies and cognitive biases such as ordering effects. This susceptibility persists across model sizes, though larger models tend to engage in more deliberate reasoning, showing a reduced sensitivity towards these errors. A notable finding indicates that LLMs, unlike humans, rarely produce the “nothing follows” response, even when it represents the accurate deduction. In a similar vein, Dasgupta et al. (2022) demonstrate that LLMs, akin to humans, exhibit content effects in predicate logic problems, i.e. their reasoning is influenced by the semantic content of the task. This effect is evident in various LLMs including ChatGPT, PaLM 2, and Chinchilla (Hoffmann et al., 2022). In particular, analyses indicate that the models’ reasoning in syllogisms is biased by the believability of the conclusion, a behavior known as belief bias (Klauer et al., 2000). Further research corroborates the influence of semantic content on the models’ reasoning. Ando et al. (2023) demonstrate a belief-bias in models like ChatGPT, RoBERTa and BART (Lewis et al., 2020) by comparing the models’ behavior on abstract, belief-consistent, and belief-inconsistent syllogisms. Their analyses also uncover a predisposition towards conversion errors and atmosphere effects, where logical quantifiers are either misinterpreted or lead to uninformed inferences, respectively (Tversky & Kahneman, 1974). Itzhak et al. (2024) further show that methods such as instruction tuning and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) may amplify cognitive biases within LLMs. Mondorf & Plank (2024) delve into the inferential strategies of open-access LLMs in problems of propositional logic. Comprehensive manual evaluations of the models’ rationales reveal that LLMs utilize inferential strategies similar to those employed by human reasoners. Their evaluations further underline difficulties of LLMs with logical negations and a vulnerability to logical fallacies commonly observed in human reasoning. Similarly, McKenzie et al. (2023) indicate that LLMs tend to replicate human-like logical errors when engaging with the logical principle of modus tollens, suggesting an imitation of flawed reasoning patterns from their training data. Notably, this trend becomes more pronounced with increasing model size, a phenomenon denoted as inverse scaling.
@@ Appendix B Further Details on Evaluation Methods
FAC2E Use capability-specific instructions to elicit intermediate structured reasoning steps (crystallized and fluid reasoning). Evaluate each step separately, using ground truth responses and the BARTScore-Recall metric. Wang et al. (2024a)